The thickness of soap film (n=1.46) is 300nm.what wavelength of light in the visible region can be used to constructively reflect off the film?

 
 
 

Grade:12

1 Answers

Arun
5 years ago
For constructive interference , 
where  is refractive index of soap film,  wavelength of light sectrum and t is thickness of soap film.
Here, t = 300nm ,  = n = 1.46 
so, 1.46 × 300 = (2n + 1)λ/4 
=> 1.46 × 300 × 4 = (2n + 1)λ
=> 1.46 × 1200/(2n + 1) = λ
for n = 1 , λ = 1.46 × 1200/3 = 1.46 × 400
λ = 5.84 × 100 = 584 nm 
for n = 2, λ = 1.46 × 1200/5 = 1.46 × 240 
λ = 350.4 nm 
but we know, visible region is 390 - 700 nm
so, λ = 584 nm ✓ and λ ≠ 350.4 nm 
hence, answer is 584 nm